Dublin University Rifle Club

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Don Baker, 30 Jellicoe Drive Belfast BT15 3LA Northern Ireland.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Don,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

Misé le meas, Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Rhona Barry, 36 Park Avenue Sandymount Dublin 4.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Rhona,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Shane Bastible, Oristown Navan Co Meath.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Shane,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Brendan Byrne, Golden Fort Grange Con Co Wicklow.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Brendan,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

John Byrne, 100 Ballinteer Close Ballinteer Dublin 16.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

John,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Robert Clarke, 113 Glebe Road Randalstown Co Antrim BT41 3DT Northern Ireland.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Robert,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Tony Cogan,y 35 Heatherview Park Aylesbury Estate Dublin 24.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Tony,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Rowan Conway, 25 Heathermount Crescent Comber Co Down BT23 5HW Northern Ireland

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Rowan,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Dave Cullen, 37 Main Street Rathdrum Co Wicklow.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Dave,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Gary Duff, 26 Grangewood Rd Belfast Co Down BT16 1GW Northern Ireland.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Gary,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

John Fitzmaurice, 50 Cherry Grove Sallins Avenue Naas Co Kildare.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

John,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Ivor Fletcher, 127 Belfast Road Antrim BT41 2AZ Northern Ireland.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Ivor,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Billy Forsyth, 143 Corke Abbey Woodbrook Bray Co Wicklow.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Billy,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Gerry Haskins, 375 Charlemont Griffith Avenue Dublin 9

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Gerry,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Robert&Michelle Heaney, 33 Brownlow Street Comber Co Down Northern Ireland

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Robert&Michelle,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Ray&Collette Kane, Bungalow 6 Casement Aerodrome Baldonnel Dublin 22.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Ray&Collette,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Alan Lewis, 6 Cherryvale Drive Glengormley Newtownabbey BT36 7UK Northern Ireland.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Alan,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Martin Mace, 29 Rossdowney Road Londonderry BT47 1PB Northern Ireland.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Martin,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Pat O’Brien, Golden Hill Manor Kilbride Blessington Co Wicklow.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Pat,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Kealan Symes, Ballard Cottage Ballinaclash Rathdrum Co Wicklow.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Kealan,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

J. K. Walsh, Carjon Lodge Cross Chapel Blessington Co Wicklow.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

J. K.,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Marty Walsh, 49 Castleknock Way Laurel Lodge Dublin 15.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Marty,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Sandy&Debbie Wilson, 9 Abbey Court Monkstown Co Dublin.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Sandy&Debbie,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Paddy Ashe, 91 Marian Park Tralee Co Kerry.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Paddy,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Kieran Barry, Coolmakee Fermoy Co Cork.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Kieran,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Leslie Fagan, 25 Dromheath Ladyswell Dublin 15.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Leslie,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

R. R. Green, 21 Magherana Park Wearingstown Co Armagh BT66 7QR Northern Ireland.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

R. R.,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Albert Jordan, Bahana Farm Enniskerry Co Wicklow.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Albert,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Tom Kelly, 20 St Laurence’s Road Chapelizod Dublin 20.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Tom,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Simon Kennedy, 28, The Rise Malahide Co Dublin.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Simon,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Ian Marnane, 18 Glasnevin Drive Glasnevin Dublin 11.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Ian,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Sean Marnane, 18 Glasnevin Drive Glasnevin Dublin 11.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Sean,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Robert McVeigh, 86 Gilford Road Lurgan Co Armagh BT66 7EB Northern Ireland.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Robert,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

John Moran, Bishops Court Upper Kill Co Kildare.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

John,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Denise O’Brien, 3 Castledowns Rd. Malahide Co. Dublin.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Denise,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Brendan O’Farrell, 14 Marian Park Rathfarnham Dublin 14.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Brendan,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Norman Smith, 12 Kingswood Park Belfast BT5 7EZ Northern Ireland.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Norman,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Terry Wearen, Duleek House Duleek Co Meath.

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

Terry,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600. Prize-giving will take place after this on the range.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

DURC club rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use if necessary – please contact me for further details on this.